OpenStudy (anonymous):

In Jim's class, 5 times as many students take the bus as those who walk. A total of 24 students walk or take the bus. How many more students take the bus than walk to school? 5, 6, 16, 20 This for my 4th grader so please explain so I can show her.

Directrix (directrix):

Let W represent the number of students who walk Because 5 times as many students take the bus than those who whose, let the bus riders be represented by 5W The sum of the number of walkers and the number of bus riders is 24. W + B = 24 but B = 5W So, W + B = 24 becomes: W + 5W = 24 Adding like terms, 6W = 24 Dividing both sides by 4, W = 4 so four students walk to school That means the number of bus riders is 5 times 4 = 20. Bus riders minus walkers = 20 - 4 = ? @Olivia102003